mass of solid iron cube of side 4cm is to be determined of the four spring balances available , the one best suited for this purpose would have 

Grade:9

1 Answers

Arun
5 years ago
Given:
side of solid iron cube=s= 4cm
density of iron= d= 7.8g/ cm3
volume of cube = s*s*s
=4×4×4
=64cm3
As we know,
density= mass/ volume 
mass= density x volume 
=7.8 *64
=499.2 gm
As the mass is 499.2 gm , the spring balance range should be 0-1000g
Least count is the smallest measurement we can measure.
As mass to be measured is in decimal value.Its better to select the least count as 10g
 
 
Spring balance with 0-1000g and least count 10g